OpenStudy (anonymous):

I need some help with math - A pair of equations is shown below: y = 3x − 5 y = 6x − 8 Part A: Explain how you will solve the pair of equations by substitution or elimination. Show all the steps and write the solution. (5 points) Part B: If the two equations are graphed, at what point will the lines representing the two equations intersect? Explain your answer

OpenStudy (anonymous):

Substitution Method Step 1: equate the two functions \[3x -5 = 6x - 8\] Step 2: isolate the variable (in our case, it's x) \[x = 1\] The two lines would intersect at point (1, -2).

OpenStudy (anonymous):

can u explain how to get how the lines would intersect @Cardinal_Carlo

OpenStudy (anonymous):

@Cardinal_Carlo like how did u find where they intersect

OpenStudy (anonymous):

@Outh-1 can u help me

OpenStudy (anonymous):

You can find the point in one of two ways: by graphing, or by doing substitution twice.

OpenStudy (anonymous):

@Outh-1 how do u do it by graphing

OpenStudy (anonymous):

Graph both lines. Try to draw it out if you can.

OpenStudy (anonymous):

So for your equations' intersection, part of the answer is in part A which we found to be x = 1, but we still need its corresponding y-value to get the intersecting point. So: \[x = 1\] \[y = f(1) = 3(1)-5 = 6(1) - 8 = -2\] Now we have both x and y values, and our point of intersection is (1, -2)

OpenStudy (anonymous):

is it supposed to come up as y = f(1) = 3(1) - 5 = 6(1) - 8 = -2 cause i dont really understand that. where do all the ones come from? sorry im really bad at math.

OpenStudy (anonymous):

@Cardinal_Carlo ^^^

OpenStudy (anonymous):

and where did the f come from @Cardinal_Carlo

OpenStudy (anonymous):

Okay, so to make things simple: Step 1 Take the solution we got from part A. Note: Solution value is synonymous with intersection value. \[x = 1\] This x-value is our solution for part A and happens to also be the intersecting x-coordinate. Step 2 Plug this value in to either equation above like so: \[y = 3(1) - 5\] and we'll get a y-value of: \[y = -2\] So now that we have both x and y values, we have the intersecting coordinates (1, -2). P.S. Never mind the 'f' if you haven't come across functions yet in your class.

OpenStudy (anonymous):

@Cardinal_Carlo thanks so much :)

OpenStudy (anonymous):

You're welcome